# Bike-Bench: A Bicycle Design Benchmark for Generative Models with Objectives and Constraints

### v1

@@ Abstract
We introduce Bike-Bench, an engineering design benchmark for evaluating generative models on problems with multiple real-world objectives and constraints. As generative AI’s reach continues to grow, evaluating its capability to understand physical laws, human guidelines, and hard constraints grows increasingly important. Engineering product design lies at the intersection of these difficult tasks, providing new challenges for AI capabilities. Bike-Bench evaluates AI models’ capability to generate designs that not only resemble the dataset, but meet specific performance objectives and constraints. To do so, Bike-Bench quantifies a variety of human-centered and multiphysics performance characteristics, such as aerodynamics, ergonomics, structural mechanics, human-rated usability, and similarity to subjective text or image prompts. Supporting the benchmark are several datasets of simulation results, a dataset of 10K human-rated bicycle assessments, and a synthetically-generated dataset of 1.4M designs, each with a parametric, CAD/XML, SVG, and PNG representation. Bike-Bench is uniquely configured to evaluate tabular generative models, LLMs, design optimization, and hybrid algorithms side-by-side. Our experiments indicate that LLMs and tabular generative models fall short of optimization and optimization-augmented generative models in both validity and optimality scores, suggesting significant room for improvement. We hope Bike-Bench, a first-of-its-kind benchmark, will help catalyze progress in generative AI for constrained multi-objective engineering design problems. Code, data, and other resources are published at decode.mit.edu/projects/bikebench/.
@@ 1 Introduction
Generative AI has recently captured widespread attention for its general-purpose problem-solving capabilities [1, 2, 3]. Despite a wealth of exploratory work [4, 5, 6, 7], generative AI has not seen widespread adoption [8] in the trillion-dollar engineering design industry [9, 10]. Engineering design can be generally characterized as the methodical decision-making process needed for the physical realization of products, systems, or other real-world solutions [11, 12, 13]. Many of engineering design’s requisite skills pose significant challenges for current Generative AI models [14, 15], including exact constraint satisfaction, adherence to quantitative and qualitative design guidelines, and intimate knowledge of multidisciplinary physical laws. To name explicit examples, in ship hull design, generative models have been found to extensively violate geometric, performance, and safety constraints, sometimes over $95\%$ of the time [16]. In large language model (LLM) benchmarks, models fail to extract precise design regulations from engineering standards [17]. Finally, in free-form structural design, generative models fall far short of optimization algorithms [18] due to their inability to learn generalizable physics rules. To excel in engineering design, generative models must overcome their imprecise constraint satisfaction, their struggle to understand both objective and subjective design guidelines, and their ignorance of physical laws. To help the community assess progression, we introduce a benchmark (Figure 1) focused on bicycle design, a problem that prominently features each of these challenges.
@@
Figure 1: Bike-Bench is a set of evaluators and metrics, built to benchmark generative models’ capability to synthesize parametric bicycle designs satisfying a variety of objectives and constraints.
@@
As a constrained, multiphysics-guided, and human-centered design benchmark, Bike-Bench fills several important voids in the existing space of generative modeling benchmarks. Unlike many image, sketch, or 3D model-based benchmarks [19, 20] used for design, Bike-Bench evaluates exclusively parametric designs. This forces benchmarked models to synthesize designs with an exact mapping to a Computer-Aided-Design (CAD) file, essentially guaranteeing a precise, ready-to-manufacture bicycle model, rather than a more abstract image, sketch, or point cloud with less-clear downstream value. Bike-Bench also differs from CAD datasets, which rarely have features to evaluate multiphysics objectives, hard constraints, or human-centered preferences [21, 22, 23]. In contrast, Bike-Bench is comprised of 10 multidisciplinary design objectives and 15 design constraints, all of which revolve around a rich set of design evaluators. These evaluators leverage datasets of physics simulations, a geometry engine and renderer, and even a dataset of real human-sourced design assessments.
@@
Unlike classic generative modeling benchmarks, Bike-Bench is more than just an exercise in the maximization of distributional similarity. To measure the practical design capabilities of a generative model, Bike-Bench also calculates aggregate validity and optimality scores over sample sets – metrics that are frequently overlooked, despite their outsized importance in engineering design [24]. For a model to simultaneously succeed in all of Bike-Bench’s metrics, it must strategically deviate from the data manifold to improve performance and constraint satisfaction in targeted ways. Bike-Bench supports tabular models, LLMs, direct design optimization algorithms, and hybrid algorithms that transcend boundaries, providing a unique opportunity to benchmark previously disparate algorithms side-by-side. Our benchmarking results suggest that generative models, and particularly LLMs, have extensive room for growth and improvement in constrained engineering design problems. Key contributions of this paper are summarized as follows:
@@
  • We introduce a synthetic dataset of 1.4M bicycle designs represented as tabular data, SVG, PNG, and XML files for Computer-Aided-Design (CAD) software. This dataset supports a variety of design generation tasks such as text-to-cad, image-to-cad, and parametric-to-image generation.
@@
  • We introduce a dataset of 10K human-sourced ratings of bicycle designs. These ratings model subjective human assessments of designs’ usability.
@@
  • We introduce Bike-Bench, a benchmark evaluating validity, optimality, and similarity of generative-model-synthesized bicycle designs. The benchmark features 25 design evaluations spanning aerodynamics, structural mechanics, ergonomics, aesthetics, geometric feasibility, and human perception of usability.
@@
  • We benchmark a variety of design generation techniques, including a large language model (OpenAI o4-mini), multiple tabular generative models, optimization-augmented generative models, and both gradient-based and heuristic design optimization algorithms.
@@ 2 Background
We provide brief overview of engineering design benchmarks for generative models, as well as benchmarks constructed to evaluate multiple classes of algorithms. A background on data-driven bicycle design is included in Appendix A.1, as well as a motivation for the selection of bicycle design as a benchmark problem.
@@
Bike-Bench joins a limited set of engineering design benchmarks designed primarily for evaluation and comparison of generative models. DesignQA [17], for example, evaluates the ability of LLMs to answer questions about engineering standards and requirements. Engineering design is highly data-limited [25], and most datasets are not configured as benchmarks [26]. Regardless, several engineering design datasets can be configured to support performance-aware generative model training [27, 28, 29, 30]. While a few of these datasets feature numerous objectives and constraints [31, 32], most only evaluate one principal objective. Bike-Bench is the first dedicated generative modeling benchmark focused on design synthesis under numerous multidisciplinary objectives and constraints.
@@
Few benchmarks directly compare generative models to entirely different classes of algorithms, because such comparisons may superficially seem ‘unfair.’ For example, even though generative models more and more frequently compete with optimization algorithms [33], the comparison may not seem fair because optimization algorithms are allowed to call evaluators, while purely data-driven models must infer from datasets. We contend that such comparisons are rigorous, important, and particularly timely. In practice, generative AI models are regularly placed in direct competition with other classes of algorithms. Direct competition across classes of models can be easily appreciated in the historical development of deep-learning-based computer vision over algorithmic methods [34] or LLMs over classical language models [35, 36]. Benchmarks that embrace this direct competition will help practitioners understand shortcomings and capacity for growth. Such benchmarking is gaining traction [37] particularly in fields like engineering design [38], where the adoption of generative AI has faced considerable resistance, due to the strength of alternative methods [18]. Bike-Bench helps fill this need for engineering design benchmarks that accommodate different classes of algorithms.
@@ 3 Datasets
Bike-Bench introduces several new datasets, many of which are used to train the predictive models that Bike-Bench uses to evaluate many facets of bicycle design. Bike-Bench also consolidates and adapts two existing datasets: The BIKED dataset [39], comprised of 4500 human-design bikes, serves as the basis for Bike-Bench’s distribution-modeling component. Bike-Bench uses a custom 70-parameter subset of salient design features from BIKED’s full representation, which are described in Appendix A.5. The FRAMED dataset [40], focusing on structural mechanics of bike frames, features nearly 15K designs simulated using Finite Element Analysis (FEA) under multiple loading cases. FRAMED supports Bike-Bench’s structural mechanics evaluators and frame validity analysis. We refer readers to the respective papers for more details on BIKED and FRAMED.
3.1 Dataset: 1.4M Synthetically-Generated Bicycle Designs
We introduce a new dataset of 1.4 million synthetically-generated bicycle designs. For each design, we provide parametric data, images, XML files and, for convenience, CLIP embeddings of the images. We generate this dataset using a customized renderer running parallelized bare-bones instances of the BikeCAD software.
@@
Synthetic datapoints are generated using CTGAN [41], a tabular GAN model specializing in realistic synthetic data generation. Each design is procedurally mapped to to a Computer-Aided-Design (CAD) file. Though CAD file structure varies widely from software to software, BikeCAD files adopt an XML format and typically consist of several thousand key-value pairs. Our custom renderer is then invoked to render the CAD-XML files into Scalable Vector Graphic (SVG) files, which are then converted to rasterized Portable Network Graphics (PNG) files. Finally, we calculate the CLIP embedding for each generated image using OpenAI’s ‘clip-vit-base-patch32’, a ViT-B/32 model provided by HuggingFace. The intermediates (XML, SVG, PNG) are retained and included in the dataset. Bike-Bench primarily uses this dataset to train a direct embedding model from Bike-Bench’s parametric design space to the CLIP embedding space, allowing for aesthetics-based design evaluation. However, the dataset can conceivably support a variety of different predictive modeling and design generation tasks, such as text-to-CAD, image-to-CAD, parametric-to-image, etc. Using the parallelized pipeline, the full 1.4M designs took approximately four days to generate on an ordinary workstation and occupy approximately 750GB of storage, compressed. Users can easily run our pipeline to generate more data at their discretion.
3.2 Dataset: 10K Human-Sourced Bicycle Ratings
@@
We also introduce a dataset of 10,350 human ratings of perceived bicycle usability. These ratings were collected through a rigorous selection procedure. Rather than spreading ratings evenly across the 4500 designs in BIKED, we instead opted to focus on a subset of 200 designs and collect approximately 50 ratings per bike, allowing us to more easily perform statistical significance tests and calculate rater agreement. This selection process for the 200 designs aimed to achieve a uniform distribution across Bike-Bench’s 70 key parameters. To target users with cycling experience and familiarity with bike usability, only countries with at least a 35% weekly riding frequency were included, limiting eligibility to 14 countries based on a survey across 28 countries in 2022 [42]. Prolific was used to crowdsource participants for rating bicycle designs, where participants answer “Yes” or “No” to the question: “Does this bicycle look easy to use?”. Binary (yes/no) assessments were selected to avoid the flaws of continuous ratings [43]. The 200 sampled bikes were divided into four groups of 50, ensuring diversity within each group. This reduced rater fatigue and helped maintain consistent rating quality throughout the session. After the completion of the rating process, a minimum rating time threshold of 90 seconds was set to remove participants who swiped too quickly, giving little time to fully evaluate each bike design. This reduced the number of valid respondents to an average of 50.75 per group, resulting in 10,350 valid ratings. Bike-Bench configures this dataset as a regression problem (predicting the proportion of raters who would consider a bike ‘easy to use’). However, a classification dataset of consensus assessments was also extracted through statistical testing, as described in Appendix A.2.2.
@@ 3.3 Dataset: 4K Cyclist Aerodynamics Simulations
We finally introduce a new dataset comprised of 4,000 3D models of cyclists in various poses and their steady-state drag force under a $10\frac{m}{s}$ relative headwind, as evaluated by a computational fluid dynamics (CFD) simulation. The 3D cyclist models are parameterized by six anthropometric measurements sampled from an approximate model of published population statistics [44, 45, 46, 47, 48], and five parameters defining the interface points between the cyclist and the bike, which are sampled from BIKED.
@@ 4 Evaluation Criteria
In this section, we discuss Bike-Bench’s design evaluation criteria. Design problems almost always feature competing objectives. Befittingly, Bike-Bench’s many objectives and constraints are challenging to optimize in synchrony. For example, the most trivial ways to optimize performance objectives will typically involve violating constraints. In this section, we discuss how Bike-Bench’s various datasets and tools are combined to evaluate a medley of functional attributes for bicycle design.
@@
In data-driven parametric design, highly expressive parametrization schemes often have a drawback in ‘allowing’ invalid configurations. Bike-Bench’s 70 design parameters enable considerable expressive flexibility, but can be selected or generated in ways that result in a variety of geometrically-infeasible designs – a common pattern among generative models trained on BIKED [49]. These designs commonly feature overlapping or disconnected components, parts with negative dimensions, and frames that violate the triangle inequality, for example. This infeasibility must therefore be identified and evaluated through constraint checks. We have compiled a set of 12 closed-form geometric constraint checks (listed in the full list of evaluations in Appendix A.4), which is significantly expanded from a set released in BIKED [39]. We supplement these closed-form checks with a data-driven feasibility predictor trained to identify bike frames with more complex geometric issues using FRAMED’s binary classification data.
Structural Soundness
@@
The structural soundness of a bike’s frame plays an important role in its comfort, power-efficiency, and safety. Bike frames are generally preferred to be as rigid as possible, which limits dissipation of energy due to the flexure of the frame when pedaling [50]. We calculate several ‘composite’ structural performance indicators – planar, transverse, and eccentric compliance, as well as frame weight – which are all considered objectives to minimize. We also calculate planar and eccentric safety factor constraints, stating that the yield stress of the frame material must be at least 1.5 times the maximum stress incurred during planar or eccentric loads. We use a predictive model trained on FRAMED [40] to these structural attributes. Similar predictive models trained on FRAMED have been shown to closely align with real-world experimental simulations [40].
Aerodynamics
@@
Aerodynamic drag reduction is a principal consideration in competitive cycling, but is generally beneficial in all cycling settings. In general, the drag force directly incurred by the bicycle is much smaller than the drag incurred by the cyclist’s body. Thus, drag is primarily minimized through repositioning of the cyclist, which is a function of the rider’s anatomical geometry and the positioning of bicycle components that the rider interfaces with (saddle, handlebars, and pedals). Therefore, aerodynamic performance is a factor of both a bicycle and an associated rider. To quantify aerodynamics, we first calculate the interface points between the bike and the cyclist based on the rider, then call a predictive model trained on the aerodynamics data for a drag force estimate.
Ergonomics
@@
In addition to playing a significant role in aerodynamics, cyclist geometry also plays an important role in ergonomics. Examining the range of angles experienced by certain key joints during regular cycling activity [51] is a simple indicator for ergonomic fit. We develop a kinematics solver which calculates a rider’s maximum knee angle, hip angle and shoulder angle and compares them to published target values for various types of cycling [51]. This solver yields three ergonomic objective scores which are calculated as a function of the bicycle design, the cyclist’s anthropomorphic measurements, and the cyclist’s use case (road biking, mountain biking, or commuting).
Human Perception of Usability
@@
Consumers frequently use subjective criteria to evaluate products. Increasingly, products are marketed as ‘user-friendly,’ with usability transitioning from a historic perception as a ‘bonus’ feature to a requisite expectation for many modern products [52]. Additionally, manufacturers emphasize their products as ‘people-oriented,’ particularly for items that involve direct body contact or require manual operation [53]. Thus, it is important to focus not only on the performance of the bicycle but also on its perceived usability. Accounting for human-centered criteria allows insights gathered from a broad population to help align the design of the bicycles, evolving them to meet growing consumer expectations. We evaluate the perceived usability of bikes using a regression model trained on our new dataset of 10K human-sourced usability scores.
Aesthetics
@@
Bicycles are regularly used for fashion, lifestyle, and other cultural statements, elevating the bicycle from a mere transportation product to a versatile tool for nuanced individual expression [54]. As such, the bicycle design industry is heavily influenced by aesthetics, individual preferences, and customization. Since individualized design customization requires significant design throughput, customization is an area where data-driven design methods can particularly excel. Thus, generative models would ideally be able to conditionally generate designs that are customized for an individual user’s aesthetic preferences, as communicated through a reference image or a text prompt. To enable this sort of conditioning, we train a predictive model to directly estimate CLIP embeddings of parametric bicycle designs using the dataset of 1.4M synthetically-generated bike designs. By calculating CLIP embeddings for parametric designs, we can quickly evaluate their similarity to text prompts or reference images.

### v2

# BikeBench: A Bicycle Design Benchmark for Generative Models with Objectives and Constraints

@@ Abstract
We introduce BikeBench, an engineering design benchmark for evaluating generative models on problems with multiple real-world objectives and constraints. As generative AI’s reach continues to grow, evaluating its capability to understand physical laws, human guidelines, and hard constraints grows increasingly important. Engineering product design lies at the intersection of these difficult tasks, providing new challenges for AI capabilities. BikeBench evaluates AI models’ capabilities to generate bicycle designs that not only resemble the dataset, but meet specific performance objectives and constraints. To do so, BikeBench quantifies a variety of human-centered and multiphysics performance characteristics, such as aerodynamics, ergonomics, structural mechanics, human-rated usability, and similarity to subjective text or image prompts. Supporting the benchmark are several datasets of simulation results, a dataset of 10,000 human-rated bicycle assessments, and a synthetically generated dataset of 1.6M designs, each with a parametric, CAD/XML, SVG, and PNG representation. BikeBench is uniquely configured to evaluate tabular generative models, large language models (LLMs), design optimization, and hybrid algorithms side-by-side. Our experiments indicate that LLMs and tabular generative models fall short of hybrid GenAI+optimization algorithms in design quality, constraint satisfaction, and similarity scores, suggesting significant room for improvement. We hope that BikeBench, a first-of-its-kind benchmark, will help catalyze progress in generative AI for constrained multi-objective engineering design problems. We provide code, data, an interactive leaderboard, and other resources at https://github.com/Lyleregenwetter/BikeBench.
@@ 1 Introduction
Generative AI has recently captured widespread attention for its general-purpose problem-solving capabilities [1, 2, 3]. Despite a wealth of exploratory work [4, 5, 6, 7], generative AI has not seen widespread adoption [8] in the trillion-dollar engineering design industry [9, 10]. Engineering design can be generally characterized as the methodical decision-making process needed for the physical realization of products, systems, or other real-world solutions [11, 12, 13]. Many of engineering design’s requisite skills pose significant challenges for current generative AI models [14, 15], including exact constraint satisfaction, adherence to quantitative and qualitative design guidelines, and intimate knowledge of multidisciplinary physical laws. In ship hull design, for example, generative models have been found to extensively violate geometric, performance, and safety constraints, sometimes over $95\%$ of the time [16]. In large language model (LLM) benchmarks, models fail to extract precise design regulations from engineering standards [17]. Finally, in free-form structural design, generative models fall far short of optimization algorithms [18] due to their inability to learn generalizable physics rules. To maximally leverage generative AI for engineering design, practitioners need models that can precisely achieve constraints, understand both objective and subjective design guidelines, and adhere to physical laws. To help the community assess progression, we introduce a benchmark (Figure 1) focused on bicycle design, a problem that prominently features each of these challenges.
@@
Figure 1: BikeBench is a set of evaluators, metrics, and datasets—built to benchmark generative models’ capability to synthesize parametric bicycle designs satisfying a variety of objectives and constraints.
@@
As a constrained, multiphysics-guided, and human-centered design benchmark, BikeBench fills several important voids in the existing space of generative modeling benchmarks. Unlike many image, sketch, or 3D model-based benchmarks [19, 20] used for design, BikeBench evaluates exclusively parametric designs. This forces benchmarked models to synthesize designs with an exact mapping to a Computer-Aided-Design (CAD) file, essentially guaranteeing a precise, ready-to-manufacture bicycle model, rather than a more abstract image, sketch, or point cloud with less-clear downstream value. BikeBench also differs from CAD datasets, which rarely have features to evaluate multiphysics objectives, hard constraints, or human-centered preferences [21, 22, 23]. In contrast, BikeBench is composed of 10 multidisciplinary design objectives and 40 design constraints, all of which revolve around a rich set of design evaluators. These evaluators leverage datasets of physics simulations, a geometry engine and renderer, and even a dataset of real human-sourced design assessments.
@@
Unlike classic generative modeling benchmarks, BikeBench is more than just an exercise in the maximization of distributional similarity. To measure the practical design capabilities of a generative model, BikeBench also calculates design quality, constraint violation, and diversity scores over sample sets—metrics that are frequently overlooked, despite their outsized importance in engineering design [24]. For a model to simultaneously succeed in all of BikeBench’s metrics, it must strategically deviate from the data manifold to improve performance and constraint satisfaction in targeted ways. BikeBench supports tabular models, LLMs, direct design optimization algorithms, and hybrid algorithms, providing a unique opportunity to benchmark previously disparate algorithms side-by-side. Our benchmarking results suggest that hybrid GenAI+optimization algorithms are current frontrunners, while LLMs have extensive room for growth and improvement in constrained engineering design problems. Key contributions of this paper are summarized as follows:
@@
  • We introduce a synthetic dataset of 1.6M bicycle designs represented as tabular data, SVG, PNG, and XML files for Computer-Aided-Design (CAD) software. This dataset supports a variety of design generation tasks such as text-to-CAD, image-to-CAD, and parametric-to-image generation.
@@
  • We introduce a dataset of 10,000 human-sourced ratings of bicycle designs. These ratings model subjective human assessments of designs’ usability.
@@
  • We introduce BikeBench, a benchmark evaluating design quality, constraint satisfaction, similarity, and diversity of generative-model-synthesized bicycle designs. The benchmark features 50 design evaluations spanning aerodynamics, structural mechanics, ergonomics, aesthetics, geometric feasibility, and human perception of usability.
@@
  • We benchmark a variety of design generation techniques, including GPT-5-high, multiple tabular generative models, optimization-augmented generative models, and both gradient-based and heuristic design optimization algorithms.
@@ 2 Background
We provide a brief overview of engineering design benchmarks for generative models, as well as benchmarks constructed to evaluate multiple classes of algorithms. A background on data-driven bicycle design is included in Appendix B.1, as well as a motivation for the selection of bicycle design as a benchmark problem.
@@
BikeBench joins a limited set of engineering design benchmarks for the evaluation and comparison of generative models. Some benchmarks evaluate question-answering capability for engineering standards [17], simulations [25], or larger taxonomies of engineering knowledge [26], but these fall short of benchmarking design synthesis capabilities. Engineering design is highly data-scarce [27], and most datasets are not configured as design synthesis benchmarks [28]. Regardless, several engineering design datasets can be configured to support performance-aware generative model training [29, 30, 31, 32]. While a few of these datasets feature numerous objectives and constraints [33, 34], most lack standardized evaluation criteria. Dedicated benchmarks for GenAI-based design synthesis have also been introduced, but primarily support just parametric GenAI models [35]. BikeBench is the first design synthesis benchmark with numerous multidisciplinary objectives and constraints to compare many classes of algorithms against various types of GenAI.
@@
Few benchmarks directly compare generative models to entirely different classes of algorithms, because such comparisons may superficially seem ‘unfair.’ For example, even though generative models more and more frequently compete with optimization algorithms [36], the comparison may not seem fair because optimization algorithms are allowed to call evaluators, while purely data-driven models must infer from datasets. We contend that such comparisons are rigorous, important, and particularly timely. In practice, generative AI models are regularly placed in direct competition with other classes of algorithms. Direct competition across classes of models can be easily appreciated in the historical development of deep-learning-based computer vision over algorithmic methods [37] or LLMs over classical language models [38, 39]. Benchmarks that embrace this direct competition will help practitioners understand shortcomings and capacity for growth. Such benchmarking is gaining traction [40], particularly in fields like engineering design [41], where the adoption of generative AI has faced considerable resistance due to the strength of alternative methods [18]. BikeBench helps fill this need for engineering design benchmarks that accommodate different classes of algorithms.
@@ 3 Datasets
BikeBench introduces several new datasets, many of which are used to train the predictive models that BikeBench uses to evaluate many facets of bicycle design. BikeBench also consolidates and adapts two existing datasets. The BIKED dataset [42], comprised of 4,500 human-design bikes, serves as the basis for BikeBench’s distribution-modeling component. BikeBench uses a custom 64-parameter subset of salient design features from BIKED’s full representation, which are described in Appendix B.5. It also uses a more selective curation and deduplication process, resulting in 3,795 designs. The FRAMED dataset [43], focusing on structural mechanics of bike frames, features nearly 15,000 designs simulated using Finite Element Analysis (FEA) under multiple loading cases. FRAMED supports BikeBench’s structural mechanics evaluators and frame validity analysis. We refer readers to the respective papers for more details on BIKED and FRAMED.
3.1 Dataset: 1.6M Synthetically Generated Bicycle Designs
We introduce a new dataset of 1.6 million synthetically generated bicycle designs. For each design, we provide parametric data, images, XML files, and, for convenience, CLIP embeddings of the images. We generate this dataset using a customized renderer running parallelized bare-bones instances of the BikeCAD software.
@@
Half of the synthetic datapoints are generated to be ‘realistic’ using CTGAN [44], a tabular GAN model specializing in realistic synthetic data generation. The other half are generated to be ‘random’ using uniform random sampling from the dataset’s most extreme bounds. Each design is procedurally mapped to a Computer-Aided-Design (CAD) file. Though CAD file structure varies widely from software to software, BikeCAD files adopt an XML format and typically consist of several thousand key-value pairs. Our custom renderer is then invoked to render the CAD-XML files into Scalable Vector Graphic (SVG) files, which are then converted to rasterized Portable Network Graphics (PNG) images. Finally, we calculate the CLIP embedding for each generated image using OpenAI’s ‘clip-vit-base-patch32’, a ViT-B/32 model provided by HuggingFace. The intermediates (XML, SVG, PNG) are retained and included in the dataset. BikeBench primarily uses this dataset to train a direct embedding model from BikeBench’s parametric design space to the CLIP embedding space, allowing for aesthetics-based design evaluation. However, the dataset has broad utility outside of the benchmarking context. It can support a wide variety of predictive modeling and design generation tasks, such as text-to-CAD, image-to-CAD, parametric-to-image, etc. Using the parallelized pipeline, the full 1.6M designs took approximately four days to generate on an ordinary workstation and occupy approximately 750GB of storage, compressed. Users can easily run our pipeline to generate more data at their discretion.
3.2 Dataset: 10,000 Human-Sourced Bicycle Ratings
@@
We also introduce a dataset of 10,350 human ratings of perceived bicycle usability. These ratings were collected through a rigorous selection procedure. Rather than spreading ratings evenly across the 4500 designs in BIKED, we instead opted to focus on a representative subset of 200 designs and collect approximately 50 ratings per bike, allowing us to more easily perform statistical significance tests and calculate rater agreement. This selection process for the 200 designs aimed to achieve a uniform distribution across BikeBench’s 64 key parameters. To target users with cycling experience and familiarity with bike usability, only countries with at least a 35% weekly riding frequency were included, limiting eligibility to 14 countries based on a survey across 28 countries in 2022 [45]. Prolific was used to crowdsource participants. These participants answered “Yes” or “No” to the question: “Does this bicycle look easy to use?”. Binary (yes/no) assessments were selected to avoid the flaws of continuous ratings [46]. The 200 sampled bikes were divided into four groups of 50, ensuring diversity within each group. This reduced rater fatigue and helped maintain consistent rating quality throughout the session. After the completion of the rating process, a minimum rating time threshold of 90 seconds was set to remove participants who swiped too quickly, giving little time to fully evaluate each bike design. This reduced the number of valid respondents to an average of 50.75 per group, resulting in 10,350 valid ratings. BikeBench configures this dataset as a regression problem (predicting the proportion of raters who would consider a bike ‘easy to use’). However, a classification dataset of consensus assessments was also extracted through statistical testing, as described in Appendix B.2.2. Information on data collection from human subjects is included in Appendix B.2.1.
@@ 3.3 Dataset: 4K Cyclist Aerodynamics Simulations
We finally introduce a new dataset composed of 4,000 simulations of 3D cyclist models in various poses. The dataset reports steady-state drag force under a $10,\mathrm{m/s}$ relative headwind, as evaluated by a computational fluid dynamics (CFD) analysis. The 3D cyclist models are parameterized by six anthropometric measurements sampled from an approximate model of published population statistics [47, 48, 49, 50, 51], and five parameters defining the interface points between the cyclist and the bike, which are sampled from BIKED.
@@ 4 Evaluation Criteria
In this section, we discuss BikeBench’s design evaluation criteria. Design problems almost always feature competing objectives. Befittingly, BikeBench’s many objectives and constraints are challenging to optimize simultaneously. For example, the most trivial ways to optimize performance objectives will typically involve violating constraints. In this section, we discuss how BikeBench’s various datasets and tools are combined to evaluate a medley of functional attributes for bicycle design.
@@
In data-driven parametric design, highly expressive parametrization schemes often have a drawback in ‘allowing’ invalid configurations. BikeBench’s 64 design parameters enable considerable expressive flexibility, but can be selected or generated in ways that result in a variety of geometrically infeasible designs—a common pattern among generative models trained on BIKED [52]. These designs commonly feature overlapping or disconnected components, parts with negative dimensions, and frames that violate the triangle inequality, for example. This infeasibility must therefore be identified and evaluated through constraint checks. We have compiled a set of 31 closed-form geometric constraint checks (listed in full in Appendix B.4), which is significantly expanded from a set released in BIKED [42]. We supplement these closed-form checks with a data-driven feasibility predictor trained on FRAMED’s binary classification data to identify bike frames with more complex geometric issues.
Structural Soundness:
@@
The structural soundness of a bike’s frame plays an important role in its comfort, power-efficiency, and safety. Bike frames are generally preferred to be as rigid as possible, which limits dissipation of energy due to the flexure of the frame when pedaling [53]. We calculate several ‘composite’ structural performance indicators—planar, transverse, and eccentric compliance, as well as frame weight—which are all considered objectives to minimize. We also calculate planar and eccentric safety factor constraints, stating that the yield stress of the frame material must exceed the maximum stress incurred during planar or eccentric loads. We use a model trained on FRAMED [43] to predict these structural attributes. Similar predictive models trained on FRAMED have been shown to closely align with real-world experimental simulations [43].
Aerodynamics:
@@
Aerodynamic drag reduction is a principal consideration in competitive cycling, but is generally beneficial in all cycling settings. In general, the drag force directly incurred by the bicycle is much smaller than the drag incurred by the cyclist’s body. Thus, drag is primarily minimized through repositioning of the cyclist, which is a function of the rider’s anatomical geometry and the positioning of bicycle components that the rider interfaces with (saddle, handlebars, and pedals). Therefore, aerodynamic performance is a factor of both a bicycle and an associated rider. To quantify aerodynamics, we first calculate the interface points between the bike and the cyclist, then call a predictive model trained on the aerodynamics data for a drag force estimate.
Ergonomics:
@@
In addition to playing a significant role in aerodynamics, cyclist geometry also plays an important role in ergonomics. Examining the range of angles experienced by certain key joints during regular cycling activity [54] is a simple indicator for ergonomic fit. We develop a kinematics solver which calculates a rider’s maximum knee angle, hip angle, and shoulder angle and compares them to published target values for various types of cycling [54]. This solver yields six validity constraints and three ergonomic objective scores which are calculated as a function of the bicycle design, the cyclist’s anthropometric measurements, and the cyclist’s use case (road biking, mountain biking, or commuting).
Human Perception of Usability:
@@
Consumers frequently use subjective criteria to evaluate products. Increasingly, products are marketed as ‘user-friendly,’ with usability transitioning from a historic perception as a ‘bonus’ feature to a requisite expectation for many modern products [55]. Additionally, manufacturers emphasize their products as ‘people-oriented,’ particularly for items that involve direct body contact or require manual operation [56]. Thus, it is important to focus not only on the performance of the bicycle but also on its perceived usability. Accounting for human-centered criteria allows insights gathered from a broad population to help align bicycles design with growing consumer expectations. We evaluate the perceived usability of bikes using a regression model trained on our new dataset of 10,000 human-sourced usability scores.
Aesthetics:
@@
Bicycles are regularly used for fashion, lifestyle, and other cultural statements, elevating the bicycle from a mere transportation product to a versatile tool for nuanced individual expression [57]. As such, the bicycle design industry is heavily influenced by aesthetics, individual preferences, and customization. Because individualized design customization requires significant design throughput, customization is an area where data-driven design methods can particularly excel. Thus, generative models would ideally be able to conditionally generate designs that are customized for an individual user’s aesthetic preferences, as communicated through a reference image or a text prompt. To enable this sort of conditioning, we train a predictive model to directly estimate CLIP embeddings of parametric bicycle designs using our dataset of 1.6M synthetic bike designs. By calculating CLIP embeddings for parametric designs, we can quickly evaluate their similarity to text prompts or reference images.